Let X and Y have the joint pdf f(x,y) = e-x-y I(x > 0,y > 0).

a. What are the marginal pdfs of X and Y ? Are X and Y independent? Why?

b. Please calculate the cumulative distribution functions for X and Y, that is, find F(x) and F(y).

c. Let Z = max(X,Y), please compute P(Z ≤ a) = P(max(X,Y) ≤ a) for a > 0. Then compute the pdf of Z.
Answer #1

f(x,y) = exy , x>0, yo (a) Marginal pdf of X £.(a) = (f(ds = j *% d an feidy =ē ( (0+1) = ēx So, fx(x) = ēt x>0. get marginaland Y are independent} let z= max(x, y) P(z = a) = (max(x,y) ca). = P(xca, ya) = P(x sa) Plysa) (since, X = (a). F, (a) = (1-